How do I find the demand equation for p=$20 sells x=42 and p=$10 sells x=52, in the form of p=mx+b using two points in the form (x,p)?
Answer by stanbon(57347)

How do I find the demand equation for p=$20 sells x=42 and p=$10 sells x=52, in the form of p=mx+b using two points in the form (x,p)?
---
Two points: (42,20) and (52,10)
---
slope = (20-10)/(42-52) = 10/-10 = -1
----
Form: p = mx + b
20 = -1(42) + b
b = 72
------------------
Equation:
p(x) = -x+72
==================
Cheers,
Stan H.